1.) A gas with an initial pressure of 1500 torr at 145 ∘C is cooled to 0 ∘C.

2.) A gas in an aerosol can at an initial pressure of 1.40 atm at 12 ∘C is heated to 39 ∘C.
Answer #1

1)
Given:
Pi = 1500 torr
Ti = 145.0 oC
= (145.0+273) K
= 418 K
Tf = 0.0 oC
= (0.0+273) K
= 273 K

use:
Pi/Ti = Pf/Tf
1500 torr / 418.0 K = Pf / 273.0 K
Pf = 979.7 torr
Answer: 979.7 torr

2)
Given:
Pi = 1.40 atm
Ti = 12.0 oC
= (12.0+273) K
= 285 K
Tf = 39.0 oC
= (39.0+273) K
= 312 K

use:
Pi/Ti = Pf/Tf
1.40 atm / 285.0 K = Pf / 312.0 K
Pf = 1.5326 atm
Answer: 1.53 atm